Two isotopic compounds are produced in the synthesis of a new therapeutic drug, which must be separated by repeated chromatography cycles. In each cycle, the solution containing both compounds passes through a chromatography column of length

1 0
Answer #1

Take your resolution and divide by the square root of the number of cycles and then solve for N to get the theoretical plates:

1.98/sqrt of 18 a= 0.46

Use the formula

R=sqrt of N/4 *(?-1) ?=unadjusted relative retention

0.46=sqrt N/4 * ( 1.093-1)

Rearrange the formula:

(0.46(4)/.093)^2=N

N= 391.44
Plate Height Formula:

N=L/H

rearrange the formula as:

H=L/N

convert the Length of 45 cm to micrometers

H=450000micrometers/(391.44)=1149.60



Now Estimate the resolution in 7 cycles.

resolution 1/resolution 2=Sqrt of (number of cycles 1/ of cycles 2)

First, take the sqrt of 7/17= 0.641

Now, 0.641 and multiply by the resolution in 17 cycles:

0.641* 1.98

=1.26